What is the base number that when added to 6% of it gets 5141.70?
(I miss typed the amount in the previous question.)
Please show the formula.

Answer by checkley77(12844)

X+(1+.06X)=5141.70
X+1.06X=5141.70
2.06X=5141.70
X=5141.70/2.06
X=2,495.97 ANS.